An assembly line worker at Rob’s Bicycle factory adds a seat to
a bicycle at a rate of 2 seats in 11 minutes. Write a proportion relating the
number of seats s to the number of minutes m. At this rate, how long will
it take to add 16 seats? 19 seats?

Answer by Boreal(15235)

2 seats/11 minutes=16 seats/x minutes.
Can cross-multiply, so that 2x=16*11=176 and x=88 minutes.
Also note that 16 is 8 x as many seats, so the time is 8*11 minutes.
For 19 seats, this is 2/11=19/x
Cross-multiply and get 2x=209 and x=104.5 minutes
Alternatively, one can say 2 seats in 11 minutes and 1 seat in 5 1/2 minutes.
Multiply 16 by 5(1/2) and get 88
multiply 19 by 5 (1/2) and get 104.5